Suppose that in the model of Myers and Majluf (1984) the parameters take on the following values: λ = 0.60, H = 3, L = 2, I = 1.0, R = 1.25, r = 10 per cent. Use the above data and answer the following questions: (a) Does the model have a separating equilibrium, a pooling equilibrium, or both? (b) Based on your answer to question (a), what is the gain in wealth of the current shareholders of a firm that announces the issuing of new shares to finance the project of investment?



1
Expert's answer
2022-02-25T10:42:20-0500

Given:

λ = 0.60

H = 3

L = 2

I = 1.0

R = 1.25

r = 10%

Solution:

a)

Common Stock λ

0.60

preferred Stock H

3

C.P Stock L

2

Convertible bonds I

1

Straight bonds R

1.25

 

r= 10%

E+ b≤ = "(\u03bb+H+L+I)\/(\u03bb+R)"

E+ b≤= "(0.60+3+2+1)\/(10+1.25)"  = 0.528

b) Based on your answer to question 

a)

The gain in wealth is given below:

Gain= 0.52 "\u00d7" 100

Gain= 52%